OpenStudy (destinyyyy):

An employment agency specializing in temporary construction help pays heavy equipment operators $121 per day and general laborers $83 per day. If 35 people were hired and the payroll was $3741, how many heavy equipment operators were employed? How many laborers?

OpenStudy (destinyyyy):

My answer is--- x= 10225.4 and y= -10190.4 are my answers right?

OpenStudy (destinyyyy):

x= equipment operators and y= laborers

OpenStudy (destinyyyy):

@whpalmer4

OpenStudy (anonymous):

you there?

OpenStudy (destinyyyy):

Yes

OpenStudy (anonymous):

ok, lets go through the steps

OpenStudy (destinyyyy):

Guess its wrong then

OpenStudy (anonymous):

ok we have 2 equtions

OpenStudy (anonymous):

x + y = 35 121x + 83y = 3741

OpenStudy (anonymous):

are those the equations you got?

OpenStudy (destinyyyy):

Yes

OpenStudy (anonymous):

great

OpenStudy (anonymous):

isolate x in the first equation x=35-y

OpenStudy (anonymous):

plug in 35-y for x in the second equation

OpenStudy (destinyyyy):

x=8,072.6

OpenStudy (anonymous):

121(35-y) + 83y = 3741

OpenStudy (destinyyyy):

I miss calculated 121x-83x.. I wrote 30 down instead of 38

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

tell me what you get

OpenStudy (destinyyyy):

x=8072.6 and y= 8037 correct?

OpenStudy (anonymous):

nope

OpenStudy (destinyyyy):

-.-

OpenStudy (anonymous):

Do you understand how i got to 121(35-y) + 83y = 3741

OpenStudy (anonymous):

I explained it above

OpenStudy (destinyyyy):

-83(x+y)=35 121x+83y=3741 -83x-83y=-2,905 --------------- 38x = -2,905 --- ----- 30 30 x=-96.83

OpenStudy (anonymous):

why -83(x+y)=35?

OpenStudy (anonymous):

it should just be x+y=35

OpenStudy (destinyyyy):

Thats what my examples are from MyMathLab

OpenStudy (destinyyyy):

It solves that way.. I have no clue what you were showing me

OpenStudy (anonymous):

I have the answer, -83(x+y)=35 doesn't make any sense

OpenStudy (destinyyyy):

-96.83+y=35 + + y=131.83

OpenStudy (anonymous):

it says only 35 people were hired, how can there be 131.83 laborers?

OpenStudy (anonymous):

Do you want me to show you how i got the answer?

OpenStudy (destinyyyy):

Hold on.. Let me start over

OpenStudy (anonymous):

These are the equations x + y = 35 121x + 83y = 3741

OpenStudy (destinyyyy):

Yes

OpenStudy (anonymous):

isolate x in the first equation x=35-y

OpenStudy (anonymous):

plug in 35-y for x in the other equation

OpenStudy (anonymous):

121(35-y) + 83y = 3741

OpenStudy (anonymous):

solve for y

OpenStudy (destinyyyy):

x=22 and y= 13 ... Do you want me to show you how i got it?

OpenStudy (anonymous):

thats what i got, good job

OpenStudy (anonymous):

to see if these are right, plug them back into the original equations

OpenStudy (destinyyyy):

-83(x-y)=35 the -83 is meant to cancel out the 83y in the other equation

OpenStudy (destinyyyy):

Im good

OpenStudy (anonymous):

22+13 = 35

OpenStudy (anonymous):

first equation works

OpenStudy (anonymous):

121(22)+83(13)=3741

OpenStudy (destinyyyy):

I already got an A on my test, Thanks for you help. It was a calculation error on my end.

OpenStudy (anonymous):

those the right answers

OpenStudy (anonymous):

k, don't forget to medal

OpenStudy (destinyyyy):

Thanks for your help.

OpenStudy (anonymous):

no problem